# Kairos: Towards Adaptive and Generalizable Time Series Foundation Models

### v1

@@ C.1.1 Model Configurations
Layers Heads $d_{\text{model}}$ $d_{\text{ff}}$ $d_{\text{expert}}$ $P$ $K$ $S$ $Z$ $\tau$ $\eta_{b}$ $w$ $\theta_{\text{init}}$ Params
@@
$\textsc{Kairos}_{\text{base}}$ 6 8 512 2048 1408 $\{32,64,128\}$ 3 3 2 $\{0.55,0.1,0.05,0.15,0.15\}$ 0.01 128 $10000^{-2j/64}$ 50M
C.1.2 Loss Function
@@
C.2 Evaluation Details
C.2.1 Metric
@@ C.2.2 Stride
C.3 Mixture-of-Size Dynamic Patching (MoS-DP)
C.3.1 Auxiliary-Loss-Free Load Balancing
@@
We define a target load distribution $\tau=(\tau_{1},\tau_{2},\ldots,\tau_{S})$, where $\tau_{i}$ is the desired proportion of the total load for expert $i$, satisfying $\sum_{j=1}^{S}\tau_{j}=1$. Next, we update the bias term $\bm{b}_{i}$ using $L_{i}$ and $\tau$:
$\bm{b}_{i}\leftarrow\bm{b}_{i}+\eta_{b}\cdot\frac{\tau_{i}\cdot\sum_{j=1}^{S}L_{j}-L_{i}}{\sum_{j=1}^{S}L_{j}},$ (14)
@@
In practical implementation, $\bm{b}_{i}$ is updated once per batch, i.e., $L_{i}$ denotes the sum of the normalized gating values for the $i$-th expert across all sequences within the entire batch. At the end of each step, the bias term $\bm{b}_{i}$ for each expert is adjusted. This dynamic adjustment of $\bm{b}_{i}$ aims to balance the workload across the experts according to the desired distribution by influencing future top-K selections, while also steering the patches-to-expert affinity scores $s^{\prime}_{n,i}$ in subsequent batches, ensuring that the actual load distribution $L_{i}$ progressively aligns with the target distribution $\tau_{i}$, thereby promoting balanced expert utilization over time.
C.3.2 Definition of the ancestor function
In this section, we provide the formal definition for the function $\mathcal{A}(\cdot,\cdot)$ within the Dynamic Patch Fusion (DPF) module, as introduced in Section 3.2.
Consider the $n$-th coarsest patch, which is initially partitioned at the coarsest granularity, $p_{S}$. As this coarsest patch is processed by the Dynamic Patch Router, $K$ experts are activated. Let $p_{k}$ denote the minimum patch size among these activated experts. The coarsest patch is subsequently partitioned into $M_{n}=\frac{p_{S}}{p_{k}}$ non-overlapping finest patches. We denote the $m$-th such finest patch as $\bm{x}_{n,m}^{p_{k}}$.
The function $\mathcal{A}(\bm{x}_{n,m}^{p_{k}},p_{i})$ is then defined as the operation that retrieves the ancestor patch of size $p_{i}$ that contains the finest patch $\bm{x}_{n,m}^{p_{k}}$. Formally, this is expressed as:
$\mathcal{A}(\bm{x}_{n,m}^{p_{k}},p_{i})=[\bm{x}_{\lfloor\frac{(m-1)\cdot p_{k}}{p_{i}}\rfloor\frac{p_{i}}{p_{k}}+1}^{p_{k}},\dots,\bm{x}_{\lfloor\frac{(m-1)\cdot p_{k}}{p_{i}}+1\rfloor\frac{p_{i}}{p_{k}}}^{p_{k}}].$ (15)
@@ C.3.2 Definition of the ancestor function
  • Set of available patch sizes: $\{p_{1},\dots,p_{S}\}=\{32,64,128\}$.
@@
  • Number of activated experts: $K=3$.
  • Sequence context length: $T=2048$.
The sequence is partitioned using the maximum patch size $p_{S}=128$, resulting in $N=\lceil\frac{T}{p_{S}}\rceil=16$ coarsest patches. For the first coarsest patch ($n=1$), suppose the activated expert indices are {1, 3, 4}. Since experts 4 and 5 are designated as null experts, the fusion granularities are determined by the non-null activated experts, corresponding to patch sizes $p_{1}=32$ and $p_{3}=128$. The minimum patch size among these is $p_{k}=32$. Consequently, this patch is partitioned into $M_{n}=\frac{p_{S}}{p_{k}}=\frac{128}{32}=4$ non-overlapping finest patches. For the third finest patch ($m=3$), the ancestor patches are computed as follows:
  • For $i=1$ and $p_{i}=p_{1}=32$:
    $\mathcal{A}(\bm{x}_{1,3}^{32},32)=\left[\bm{x}_{\lfloor\frac{(3-1)\cdot 32}{32}\rfloor\frac{32}{32}+1}^{32},\dots,\bm{x}_{\lfloor\frac{(3-1)\cdot 32}{32}+1\rfloor\frac{32}{32}}^{32}\right]=[\bm{x}_{3}^{32}].$ (16)
    The ancestor patch is the finest patch itself.
  • For $i=2$ and $p_{i}=p_{2}=64$:
    $\mathcal{A}(\bm{x}_{1,3}^{32},64)=\left[\bm{x}_{\lfloor\frac{(3-1)\cdot 32}{64}\rfloor\frac{64}{32}+1}^{32},\dots,\bm{x}_{\lfloor\frac{(3-1)\cdot 32}{64}+1\rfloor\frac{64}{32}}^{32}\right]=[\bm{x}_{3}^{32},\bm{x}_{4}^{32}].$ (17)
    Expert 2 is not activated, so its normalized gating value, $\alpha_{n,i}$, in Equation 3 would be zero, meaning it does not contribute to the information fusion.
  • For $i=3$ and $p_{i}=p_{3}=128$:
    $\mathcal{A}(\bm{x}_{1,3}^{32},128)=\left[\bm{x}_{\lfloor\frac{(3-1)\cdot 32}{128}\rfloor\frac{128}{32}+1}^{32},\dots,\bm{x}_{\lfloor\frac{(3-1)\cdot 32}{128}+1\rfloor\frac{128}{32}}^{32}\right]=[\bm{x}_{1}^{32},\bm{x}_{2}^{32},\bm{x}_{3}^{32},\bm{x}_{4}^{32}].$ (18)
This mechanism, enabled by the function $\mathcal{A}(\cdot,\cdot)$, allows for the dynamic selection of different granularities for information fusion, tailored to the information density of each coarsest patch. Consequently, it effectively addresses the challenge of time series heterogeneity inherent in TSFMs.
C.4 Instance-Adaptive Rotary Position Embedding (IARoPE)
@@
In this section, we provide a more detailed description of the IARoPE implementation discussed in Section 3.3. We first introduce the principles of Rotary Position Embedding (RoPE) (Su et al., 2024) in detail, and then provide corresponding supplementary information on the IARoPE implementation.
C.4.1 Details of RoPE
@@
The core idea of RoPE is to rotate segments of the query and key vectors by an angle proportional according to their absolute position in the sequence. This allows the model to discern relative positions through the geometry of these rotations, without needing explicit calculation of relative distances.
@@
Specifically, RoPE operates on vectors of an even dimension, denoted as $D_{h}$. For a vector $\bm{e}$ (representing a query $\mathbf{q}$ or a key $\mathbf{k}$ in self-attention) at position $m$, it is transformed by a rotation matrix $\mathbf{R}_{m}$. This matrix is block-diagonal, composed of $D_{h}/2$ individual $2\times 2$ rotation blocks. Each block $\mathbf{R}_{m,j}$ acts on a pair of dimensions $(e_{2j},e_{2j+1})$ of the vector:
@@
As introduced in Section 3.3, RoPE applies a rotation to input vector. This rotation is applied to pairs of dimensions $(e_{2j},e_{2j+1})$, and can be concisely expressed using complex form:
$f_{\text{RoPE}}(\bm{e},m)=(e_{2j}+ie_{2j+1})e^{im\theta_{j}},$ (19)
@@
where $\theta_{j}$ is the angular frequency. This multiplication by $e^{im\theta_{j}}$ corresponds to a rotation in the complex plane. For the real-valued components $e_{2j}$ and $e_{2j+1}$, this operation is equivalent to applying the following $2\times 2$ rotation matrix $\mathbf{R}_{m,j}$:
$\mathbf{R}_{m,j}=\begin{bmatrix}\cos(m\theta_{j})&-\sin(m\theta_{j})\\ \sin(m\theta_{j})&\cos(m\theta_{j})\end{bmatrix}$, (20)
@@
here, $m$ is the absolute position of the token. The term $\theta_{j}$ represents a predefined angular frequency for the $j$-th pair of dimensions ($j\in[0,D_{h}/2-1]$), typically defined as $\theta_{j}=b^{-2j/D_{h}}$ where $b=10000$ in RoPE (Su et al., 2024) original setting.
@@
C.4.2 Details of IARoPE
The initial RoPE frequencies $\theta_{\text{init},j}$ typically exhibit a wide numerical range. For instance, in our setting $\theta_{\text{init},0}=1.0\ (j=0)$, while $\theta_{\text{init},31}\approx 1.33\times 10^{-4}$ ($j=31$, corresponding to $D_{h}/2-1$ and $D_{h}=64$). Directly applying affine modulation to these values could lead to numerical instability or disproportionate adjustments due to the vast scale differences.
To address this and ensure stable and effective modulation across the entire range of base frequencies, our IARoPE performs the adaptation in log-space. The layer-specific modulation parameters, $\gamma^{(l)}_{j}$ and $\beta^{(l)}_{j}$, predicted by the Multilayer Perceptron (MLP) for layer $l$ following Algorithm 1, are applied to the log-transformed base frequencies $\log\theta_{\text{init},j}$ via an element-wise affine transformation as
$\log\theta^{\prime(l)}_{j}=\gamma^{(l)}_{j}\odot\log\theta_{\text{init},j}+\beta^{(l)}_{j}$, (25)
where $\log\theta^{\prime(l)}_{j}$ represents the modulated log-frequencies for layer $l$ and dimension pair $j$. Then, these modulated log-frequencies are transformed back to their original scale by exponentiation to obtain the adaptive rotation frequencies $\theta^{\prime(l)}_{j}$ used in the IARoPE calculation for layer $l$ as
$\theta^{\prime(l)}_{j}=\exp(\log\theta^{\prime(l)}_{j})$. (26)
@@ C.4.2 Details of IARoPE
1:Time series instance $X\in\mathbb{R}^{B\times L}$, mask $M\in\{0,1\}^{B\times L}$, batch size $B$, sequence length $L$, FFT feature dimension $w$, MLP network $f_{\text{MLP}}$
2:Adaptive parameters $\gamma^{(l)}\in\mathbb{R}^{B\times D_{h}/2}$, $\beta^{(l)}\in\mathbb{R}^{B\times D_{h}/2}$
@@
9:return $\gamma^{(l)},\beta^{(l)}$
Appendix D Calculation of information density
$H_{SE}(\mathbf{w}_{i})=-\sum_{k=0}^{M-1}p_{i}[k]\log_{2}(p_{i}[k]).$ (28)
$\mu_{SE}=\frac{1}{N}\sum_{i=1}^{N}H_{SE}(\mathbf{w}_{i}).$ (29)

### v2

@@ D.1.1 Model Configurations
Layers Heads $d_{\text{model}}$ $d_{\text{ff}}$ $d_{\text{expert}}$ $P$ $K$ $G$ $Z$ $\tau$ $\eta_{b}$ $w$ $\theta$ Params
@@
$\textsc{Kairos}_{\text{base}}$ 6 8 512 2048 1408 $\{32,64,128,256\}$ 4 4 2 $\{0.55,0.1,0.05,0.15,0.15\}$ 0.01 128 $10000^{-2j/64}$ 53M
D.1.2 Loss Function
@@
D.2 Evaluation Details
D.2.1 Metric
@@ D.2.2 Stride
D.3 Mixture-of-Size Encoder
D.3.1 Auxiliary-Loss-Free Load Balancing
@@
We define a target load distribution $\tau=(\tau_{1},\tau_{2},\ldots,\tau_{S})$, where $\tau_{i}$ is the desired proportion of the total load for expert $i$, satisfying $\sum_{j=1}^{S}\tau_{j}=1$. Next, we update the bias term $\boldsymbol{b}_{i}$ using $L_{i}$ and $\tau$:
$\boldsymbol{b}_{i}\leftarrow\boldsymbol{b}_{i}+\eta_{b}\cdot\frac{\tau_{i}\cdot\sum_{j=1}^{S}L_{j}-L_{i}}{\sum_{j=1}^{S}L_{j}},$ (13)
@@
In practical implementation, $\boldsymbol{b}_{i}$ is updated once per batch, i.e., $L_{i}$ denotes the sum of the normalized gating values for the $i$-th expert across all sequences within the entire batch. At the end of each step, the bias term $\boldsymbol{b}_{i}$ for each expert is adjusted. This dynamic adjustment of $\boldsymbol{b}_{i}$ aims to balance the workload across the experts according to the desired distribution by influencing future top-K selections, while also steering the patches-to-expert affinity scores $s^{\prime}_{n,i}$ in subsequent batches, ensuring that the actual load distribution $L_{i}$ progressively aligns with the target distribution $\tau_{i}$, thereby promoting balanced expert utilization over time.
D.3.2 Definition of the Expansion Function
In this section, we provide the formal definition for the function $\mathrm{Expand}(\cdot)$ within the Mixture-of-Size Fusion stage, as introduced in Section 3.2.
Consider a single time series segment $\mathbf{x}$ of length $P$. During the Top-$K$ Granularity Routing, a set of valid experts $\mathcal{I}_{\mathrm{valid}}$ is activated. Let $P_{\min}$ denote the finest granularity among these activated experts:
$P_{\min}=\min\{P_{j}\mid j\in\mathcal{I}_{\mathrm{valid}}\}.$ (14)
Consequently, the target sequence length for alignment is defined as $M=\lceil P/P_{\min}\rceil$.
For an activated granularity level $i$ with patch size $P_{i}$, the encoding process yields a sequence of embeddings $\boldsymbol{E}_{i}=[\boldsymbol{e}_{i,1},\dots,\boldsymbol{e}_{i,M_{i}}]$, where $M_{i}=\lceil P/P_{i}\rceil$. Due to the nested design of patch sizes, the ratio $R_{i}=P_{i}/P_{\min}$ is strictly a positive integer.
The function $\mathrm{Expand}(\boldsymbol{E}_{i})$ maps the coarser embedding sequence to the finest temporal resolution by repeating each embedding vector $R_{i}$ times. Formally, the $k$-th element of the expanded sequence $\tilde{\boldsymbol{E}}_{i}\in\mathbb{R}^{M\times D_{h}}$ is derived from the original sequence via index mapping:
$\tilde{\boldsymbol{e}}_{i,m}=\boldsymbol{e}_{i,\lceil m/R_{i}\rceil},\quad\text{for }m=1,\dots,M.$ (15)
In vector notation, this operation can be expressed as:
$\mathrm{Expand}(\boldsymbol{E}_{i})=[\underbrace{\boldsymbol{e}_{i,1},\dots,\boldsymbol{e}_{i,1}}_{R_{i}\text{ times}},\dots,\underbrace{\boldsymbol{e}_{i,M_{i}},\dots,\boldsymbol{e}_{i,M_{i}}}_{R_{i}\text{ times}}].$ (16)
@@ D.3.2 Definition of the Expansion Function
  • Segment length: $P=128$.
  • Set of available granularity patch sizes: $\{P_{1},P_{2},P_{3}\}=\{32,64,128\}$.
@@
  • Top-$K$ selection: $K=2$.
Suppose for a specific segment $\mathbf{x}$, the router activates experts corresponding to granularity levels 1 and 3 (i.e., $\mathcal{I}_{\mathrm{valid}}=\{1,3\}$). The finest activated resolution is $P_{\min}=\min(32,128)=32$. The target aligned sequence length is $M=128/32=4$.
The expansion process for each expert is computed as follows:
  • For Expert 1 ($P_{1}=32$):
    The number of original patches is $M_{1}=128/32=4$. The embedding sequence is $\boldsymbol{E}_{1}=[\boldsymbol{e}_{1,1},\boldsymbol{e}_{1,2},\boldsymbol{e}_{1,3},\boldsymbol{e}_{1,4}]$. The repetition factor is $R_{1}=32/32=1$.
    $\tilde{\boldsymbol{E}}_{1}=\mathrm{Expand}(\boldsymbol{E}_{1})=[\boldsymbol{e}_{1,1},\boldsymbol{e}_{1,2},\boldsymbol{e}_{1,3},\boldsymbol{e}_{1,4}].$ (17)
    Since this expert operates at the finest activated resolution, the expansion is an identity mapping.
  • For Expert 3 ($P_{3}=128$):
    The number of original patches is $M_{3}=128/128=1$. The embedding sequence consists of a single vector $\boldsymbol{E}_{3}=[\boldsymbol{e}_{3,1}]$. The repetition factor is $R_{3}=128/32=4$.
    $\tilde{\boldsymbol{E}}_{3}=\mathrm{Expand}(\boldsymbol{E}_{3})=[\boldsymbol{e}_{3,1},\boldsymbol{e}_{3,1},\boldsymbol{e}_{3,1},\boldsymbol{e}_{3,1}].$ (18)
    The single coarse embedding is broadcast across all 4 time steps to align with the target resolution.
  • Hypothetical Case for Expert 2 ($P_{2}=64$):
    Although Expert 2 was not activated in this example, if it were processed, $M_{2}=128/64=2$ and $R_{2}=64/32=2$. The expansion would be:
    $\tilde{\boldsymbol{E}}_{2}=[\boldsymbol{e}_{2,1},\boldsymbol{e}_{2,1},\boldsymbol{e}_{2,2},\boldsymbol{e}_{2,2}].$ (19)
By explicitly aligning all representations to the finest resolution $P_{\min}$ via $\mathrm{Expand}(\cdot)$, Kairos ensures that the subsequent weighted summation $\bar{\boldsymbol{E}}=\sum_{i\in\mathcal{I}_{\mathrm{valid}}}\alpha_{i}\,\tilde{\boldsymbol{E}}_{i}$ is performed on element-wise compatible tensors, effectively fusing global context with local details.
D.4 Dynamic Rotary Position Embedding (DRoPE)
@@
In this section, we provide a more detailed description of the DRoPE implementation discussed in Section 3.3. We first introduce the principles of Rotary Position Embedding (RoPE) (Su et al., 2024) in detail, and then provide corresponding supplementary information on the DRoPE implementation.
D.4.1 Details of RoPE
@@
The core idea of RoPE is to rotate segments of the query and key vectors by an angle proportional according to their absolute position in the sequence. This allows the model to discern relative positions through the geometry of these rotations, without needing explicit calculation of relative distances. Throughout this subsection, we use $m$ and $n$ to denote generic token positions.
@@
Specifically, RoPE operates on vectors of an even dimension, denoted as $D_{h}$. For a vector $\boldsymbol{e}$ (representing a query $\mathbf{q}$ or a key $\mathbf{k}$ in self-attention) at position $m$, it is transformed by a rotation matrix $\mathbf{R}_{m}$. This matrix is block-diagonal, composed of $D_{h}/2$ individual $2\times 2$ rotation blocks. Each block $\mathbf{R}_{m,d}$ acts on a pair of dimensions $(e_{2d},e_{2d+1})$ of the vector:
@@
As introduced in Section 3.3, RoPE applies a rotation to input vector. This rotation is applied to pairs of dimensions $(e_{2d},e_{2d+1})$, and can be concisely expressed using complex form:
$f_{\text{RoPE}}(\boldsymbol{z},m)=(z_{2d}+iz_{2d+1})e^{im\theta_{d}},$ (20)
@@
where $\theta_{d}$ is the angular frequency. Here, $i$ denotes the imaginary unit. This multiplication by $e^{im\theta_{d}}$ corresponds to a rotation in the complex plane. For the real-valued components $z_{2d}$ and $z_{2d+1}$, this operation is equivalent to applying the following $2\times 2$ rotation matrix $\mathbf{R}_{m,d}$:
$\mathbf{R}_{m,d}=\begin{bmatrix}\cos(m\theta_{d})&-\sin(m\theta_{d})\\ \sin(m\theta_{d})&\cos(m\theta_{d})\end{bmatrix}$, (21)
@@
Here, $m$ is the absolute position of the token. The term $\theta_{d}$ represents a predefined angular frequency for the $d$-th pair of dimensions ($d\in[0,D_{h}/2-1]$), typically defined as $\theta_{d}=b^{-2d/D_{h}}$
@@
Equation 25 confirms that the interaction relies on the relative position nmn-m through the full rotation matrix $\mathbf{R}_{n-m}$. Since this matrix is block-diagonal, the total interaction is actually composed of independent rotations on disjoint 2D subspaces. The following theorem demonstrates the periodicity of the attention score within each subspace.
Theorem D.1 (Periodic dependence of RoPE attention).
Consider the $d$-th 2D subspace of the hidden state under RoPE, with angular frequency $\theta_{d}$, and let $\mathrm{atten}_{d}(m,n)$ denote this subspace’s contribution to the dot-product attention between positions $m$ and $n$. Then there exist an amplitude $A_{d}(m,n)\geq 0$ and a phase $\varphi_{d}(m,n)$, depending only on the content vectors at $m$ and $n$, such that
$\begin{split}\mathrm{atten}_{d}(m,n)&=A_{d}(m,n)\,\cos\big(\theta_{d}(m-n)+\varphi_{d}(m,n)\big)\\ &\;\propto\;\cos\big(\theta_{d}(m-n)+\varphi_{d}(m,n)\big).\end{split}$ (26)
In particular, for fixed content, $\mathrm{atten}_{d}(m,n)$ is a periodic function of the relative distance (mn)(m-n) with period $2\pi/\theta_{d}$.
Proof.
Let $q_{m,d},k_{n,d}\in\mathbb{C}$ be the complex-valued representations of the $d$-th 2D subspace for the query and key at positions $m$ and $n$, respectively, as defined above. Applying RoPE multiplies these components by phase factors $e^{im\theta_{d}}$ and $e^{in\theta_{d}}$, yielding
$q^{\prime}_{m,d}=q_{m,d}e^{im\theta_{d}},\qquad k^{\prime}_{n,d}=k_{n,d}e^{in\theta_{d}}.$ (27)
The contribution of this subspace to the dot-product attention is the real part of $q^{\prime}_{m,d}(k^{\prime}_{n,d})^{*}$:
$\mathrm{atten}_{d}(m,n)=\mathrm{Re}\!\big[q^{\prime}_{m,d}(k^{\prime}_{n,d})^{*}\big]=\mathrm{Re}\!\big[q_{m,d}k_{n,d}^{*}e^{i\theta_{d}(m-n)}\big].$ (28)
Writing $q_{m,d}k_{n,d}^{*}$ in polar form as $q_{m,d}k_{n,d}^{*}=A_{d}(m,n)e^{i\varphi_{d}(m,n)}$ with $A_{d}(m,n)\geq 0$ and $\varphi_{d}(m,n)\in\mathbb{R}$, we obtain
$\begin{split}\mathrm{atten}_{d}(m,n)&=\mathrm{Re}\!\big[A_{d}(m,n)e^{i(\theta_{d}(m-n)+\varphi_{d}(m,n))}\big]\\ &=A_{d}(m,n)\,\cos\big(\theta_{d}(m-n)+\varphi_{d}(m,n)\big),\end{split}$ (29)
which shows that $\mathrm{atten}_{d}(m,n)$ is proportional to a cosine function of the relative distance (mn)(m-n), with angular frequency $\theta_{d}$. The periodicity with period $2\pi/\theta_{d}$ follows immediately from the periodicity of the cosine function. ∎
D.4.2 Details of DRoPE
The initial RoPE frequencies $\theta_{d}$ typically exhibit a wide numerical range. For instance, in our setting $\theta_{0}=1.0\ (d=0)$, while $\theta_{31}\approx 1.33\times 10^{-4}$ ($d=31$, corresponding to $D_{h}/2-1$ and $D_{h}=64$). Directly applying affine modulation to these values could lead to numerical instability or disproportionate adjustments due to the vast scale differences.
To address this and ensure stable and effective modulation across the entire range of base frequencies, our DRoPE performs the adaptation in log-space. The layer-specific modulation parameters, $\gamma^{(l)}_{d}$ and $\beta^{(l)}_{d}$, predicted by the Multilayer Perceptron (MLP) for layer $l$ following Algorithm 1, are applied to the log-transformed base frequencies $\log\theta_{d}$ via an element-wise affine transformation as
$\log\theta^{\prime(l)}_{\text{inst},d}=\gamma^{(l)}_{d}\odot\log\theta_{d}+\beta^{(l)}_{d}$, (30)
where $\log\theta^{\prime(l)}_{\text{inst},d}$ represents the modulated log-frequencies for layer $l$ and dimension pair $d$. Then, these modulated log-frequencies are transformed back to their original scale by exponentiation to obtain the adaptive rotation frequencies $\theta^{\prime(l)}_{\text{inst},d}$ used in the DRoPE calculation for layer $l$ as
$\theta^{\prime(l)}_{\text{inst},d}=\exp(\log\theta^{\prime(l)}_{\text{inst},d})$. (31)
@@ D.4.2 Details of DRoPE
0: Time series instance $X\in\mathbb{R}^{B\times L}$, mask $M\in\{0,1\}^{B\times L}$, batch size $B$, sequence length $L$, FFT feature dimension $w$, MLP network $f_{\text{MLP}}$
0: Adaptive parameters $\gamma^{(l)}\in\mathbb{R}^{B\times D_{h}/2}$, $\beta^{(l)}\in\mathbb{R}^{B\times D_{h}/2}$
@@
7:return $\gamma^{(l)},\beta^{(l)}$
D.4.3 Interpretation of DRoPE
From the theorem D.1, the base RoPE parameters $\{\theta_{d}\}_{d=0}^{D_{h}/2-1}$ define a fixed bank of sinusoidal kernels over relative positions. Each $\theta_{d}$ controls the frequency of one cosine kernel. This fixed frequencies works reasonably well when sequences share a homogeneous notion of temporal scale, but it can become suboptimal for a TSFM that must handle highly heterogeneous sampling intervals and spectral patterns across domains.
DRoPE addresses this by making these frequencies instance-adaptive. The FFT-based module in Algorithm 1 extracts simple spectral statistics $X^{\prime}_{\text{FFT}}$ from each input sequence and maps them, via an MLP, to instance- and layer-specific parameters $\gamma^{(l)}_{d}$ and $\beta^{(l)}_{d}$. These parameters modulate the base frequencies in log-space:
$\log\theta^{\prime(l)}_{\text{inst},d}=\gamma^{(l)}_{d}\odot\log\theta_{d}+\beta^{(l)}_{d},\quad\theta^{\prime(l)}_{\text{inst},d}=\exp\big(\log\theta^{\prime(l)}_{\text{inst},d}\big)$, (32)
yielding an adapted frequency grid $\{\theta^{\prime(l)}_{d}\}$ for each layer and sequence.
Intuitively, this log-space affine transformation stretches or compresses the original RoPE frequency grid in a sequence-dependent way, while preserving the relative-position nature of RoPE. The resulting $\theta^{\prime(l)}_{\text{inst},d}$ still define sinusoidal kernels over relative lags, but their effective frequencies are now gently steered by the spectrum of the current input. Crucially, DRoPE does not aim to recover a single true period for each series; real-world time series are often multi-periodic and non-stationary. Instead, we interpret $\{\theta^{\prime(l)}_{\text{inst},d}\}$ as a sequence-dependent frequency profile that shapes how attention depends on relative lags, providing a more flexible and data-driven positional bias than fixed RoPE.